Find the general solution of the differential equations taking into account the initial conditions using the parameter variation method:

. y + 4y = t y(0) = y(0) = 0 et y(0) = 1 yiv + 2y + y = 3t+4 ; y(0) = y(0) = 0 et y(0) = y(0) = 1 y – 3y + 2y =t
Answer #1

y+ys y0=1 1 De 33²+20 (alten) (uten) D(2000-2) -Xe+ old-sty ly 3 yo) = yloso, Auxiliary egh it m+4m=0 7 m/m²+4)=0 mlm tai]